OpenStudy (anonymous):

Medal for help :) -2c=-18 A. c=-9*** My Answer B. c=-36 C. c=9 D. c=36

OpenStudy (anonymous):

There's a problem here, look at your signs.

OpenStudy (anonymous):

What happens when you divide a negative by a negative?

OpenStudy (anonymous):

you get a negative

OpenStudy (anonymous):

No you get a positive, it's like multiplication when you multiply 2 negatives you get a positive. - x - = + -/- = +

jimthompson5910 (jim_thompson5910):

Also, you can check your answer -2c=-18 -2(-9)=-18 ... replace c with -9 (since you claim c = -9) 18 = -18 .. as iambatman is saying, negative times negative = positive The equation is false. So c = -9 is not the solution to -2c=-18

OpenStudy (anonymous):

Ok I got C. c=9 as my answer

jimthompson5910 (jim_thompson5910):

much better, let's check that -2c=-18 -2*(9)=-18 ... plug in c = 9 -18 = -18 ... it works now